OpenStudy (anonymous):

find the value of r so that the line that passes through each pair of points has the given slope. (8,-2), (r,-6), m=-4

OpenStudy (anonymous):

simply use the formula to find the slope

OpenStudy (anonymous):

thus given (-6 - (-2) )/(r-8 )= -4 -4/(r-8) = -4 or r-8=1 or r=1+8=9

OpenStudy (anonymous):

I see thanks!